# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

### v1

@@ 4.2 Segmentation Datasets
Table 3: Detection performance on the LM-O and YCB-V datasets. OLN* is trained alongside the matching head whereas OLN uses fixed modules. †means the model is trained on OWID and real images. Our running time of YCB-V is based on the scene 48. ()($\cdot$)( ⋅ ) represents the number of instances.
@@
Table 4: Novel instance segmentation results on the seven core datasets of the BOP benchmark. We utilize Average Precision (AP) to compare these methods. ()($\cdot$)( ⋅ ) includes the number of instances.
@@
NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
4.3 Benchmarking Results
Detection The high-resolution dataset results are presented in Table 1. Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories. Our basic method surpasses the top baseline by 17.7 AP. Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP. For the RoboTools dataset, detailed in Table 2, our method outperforms the state-of-the-art VoxDet by 46.2 AP. Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance. Our approach achieves over 60 AP on these two fully real datasets.
Segmentation CNOS [64] and SAM-6D [49] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2. In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline. The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table 4. Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
Qualitative result We present detection results for both simple and complex scenes in Fig. 3. Additional detection and segmentation examples are available in the Appendix.
Runtime The tables of detection results include the runtime of our method. Despite Grounded-SAM reducing the number of proposals, our approach remains slower compared to end-to-end detectors. This is primarily due to the process involved in obtaining proposal embeddings. The inclusion of adapters does not impede the overall framework. Future research could focus on developing a more efficient method for robotic applications. Please see the appendix for segmentation runtime.
4.4 Ablation Study
Grounded SAM (GS) vs SAM On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5. The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone. Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
FFA vs $cls$ token To compare these two types of embedding generation, we present the results in Table 5. Our weight adapter enhances both two types of embeddings. Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter.
@@ 4.4 Ablation Study
Dinov2 backbones with adapter Our Weight Adapter is compatible with various backbones of Dinov2. Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements. Details of this comparison are provided in Table 6. Please refer to the Appendix for additional ablation studies.
5 Discussions
However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires a longer running time and greater computational resources compared to end-to-end detectors. Additionally, Grounding DINO may produce large bounding boxes that encompass multiple objects, leading to under-segmentation and detection failures. Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
In this work, instances are represented by $K$ template embeddings. For future research, we propose that each instance should be characterized by a single, distinctive embedding that acts as its identifier to enable one-shot detection. This approach will allow the detector to identify and locate a target instance within a query image using just one template image. Additionally, developing an efficient method for this process will be crucial for its application in robotics.
Acknowledgments

### v2

@@ 4.2 Segmentation Datasets
Table 3: Detection performance on the LM-O and YCB-V datasets. OLN* is trained alongside the matching head whereas OLN uses fixed modules. †means the model is trained on OWID and real images. ()($\cdot$)( ⋅ ) represents the number of instances.
@@
Table 4: Novel instance segmentation results on the seven core datasets of the BOP benchmark. We utilize Average Precision (AP) to compare these methods. ()($\cdot$)( ⋅ ) includes the number of instances. SAM6D (RGBD) exclusively utilizes RGB-D images, whereas other models employ only RGB images.
@@
NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
@@
Figure 5: Comparison of segmentation results using CNOS, SAM6D, and our NIDS-Net on the YCB-V dataset. CNOS and SAM6D may misclassify some background regions or object parts as objects due to their reliance on SAM for proposal generation. NIDS-Net addresses this limitation with Grounded-SAM.
4.3 Benchmarking Results
Detection. The high-resolution dataset results are presented in Table 1. Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories. Our basic method surpasses the top baseline by 17.7 AP. Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP. For the RoboTools dataset, detailed in Table 2, our method outperforms the state-of-the-art VoxDet by 46.2 AP. Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance. Our approach achieves over 60 AP on these two fully real datasets.
Segmentation. CNOS [66] and SAM-6D [53] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2. In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline. The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table 4. Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
Qualitative Result. We present detection results for both simple and complex scenes in Fig. 4. Fig. 5 displays a visual comparison of the segmentation results. Additional examples are available in the appendix.
Runtime. The tables of detection results include the runtime of our method. Grounded-SAM substantially decreases the number of proposals compared to SAM, thereby accelerating the entire process. Please see the appendix for segmentation runtime.
4.4 Ablation Study
Grounded SAM (GS) vs SAM. On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5. The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone. Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
FFA vs $cls$ token. To compare these two types of embedding generation, we present the results in Table 5. Our weight adapter enhances both two types of embeddings. Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter. Additional ablation studies are provided in the appendix.
5 Discussions
However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors. Additionally, Grounding DINO may produce large bounding boxes that contain multiple objects, leading to detection failures. Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
In this work, instances are represented by $K$ template embeddings. For future research, we will explore using a single, distinctive embedding for each instance that acts as its identifier to enable one-shot detection. This approach will allow the detector to identify and locate a target instance within a query image using just one template image. Additionally, developing a computationally efficient method for this process will be crucial for its application in robotics.
Acknowledgments
Supplementary Material
6 Training Details
Detection. For detection datasets, the weight adapter is trained with a batch size of 1024, while the CLIP-Adapter is trained with a batch size of 512 to enhance performance. Both adapters are trained for the same number of epochs, as detailed in Table 6. To utilize Grounding DINO, a box threshold of 0.15 is set for the high-resolution and YCB-V datasets, and 0.10 for other datasets.
Table 6: The training epochs of different detection datasets.
High-resolution RoboTools LM-O YCB-V
Both Adapters 40 80 40 40
Segmentation. We combine all instances from seven core datasets of the BOP benchmark. We train both adapters with a batch size of 1344 (32 instances $\times$ 42 templates per instance) for 500 epochs. For Grounding DINO, a box threshold of 0.10 is set for all datasets.
7 More Ablation Study
FFA vs $cls$ token and Grounded SAM (GS) vs SAM in Detection Tasks. The detection capabilities of FFA and the $cls$ token are demonstrated using the high-resolution dataset, as presented in Table 9. For instance embedding generation, FFA achieves significantly greater enhancements compared to the $cls$ token. Furthermore, Table 9 illustrates that GS is effective in generating precise object proposals.
@@ 7 More Ablation Study
Table 8: Detection results on the High-resolution real world detection dataset using various SAM variants.
SAM Variant AP AP50 AP75 Time (sec)
Mobile SAM (Tiny ViT) [97] 54.5 70.5 62.0 6.80
Mobile SAM (Tiny ViT) [97] + WA 58.2 75.0 66.5 6.73
HQ-SAM (ViT-H) [41] 59.8 71.5 65.7 7.38
HQ-SAM (ViT H) [41] + WA 63.8 76.2 70.5 7.31
SAM (ViT-H) (Ours) [47] 59.3 71.1 65.1 6.92
SAM (ViT-H) + WA (Ours) [47] 63.9 76.6 70.6 6.78
Table 9: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
Proposal Embedding AP AP50 AP75
SAM $cls$ token 41.6 49.1 46.0
GS $cls$ token 54.9 65.4 60.1
GS FFA 59.3 71.1 65.1
Image encoder. Given the same object proposals with GS, we evaluate the FFA embeddings from different image encoders on the High-resolution dataset [81]. As illustrated in Table 10, DINOv2 exhibits superior performance attributable to its robust visual features. Fig. 6 presents the visual results of various image encoders.
Table 10: The instance embeddings of different image encoders.
Image Encoder Feature Dimension AP AP50 AP75
SAM (Vit-L) [47] 256 0.9 1.1 0.9
SAM (Vit-L) [47] + WA 256 0.7 0.8 0.7
DeiT III [86] (Vit-L) 784 2.9 3.6 3.1
DeiT III [86] (Vit-L) + WA 784 2.8 3.4 3.0
CLIP (Vit-L) [75] 1024 17.9 21.1 18.9
CLIP (Vit-L) [75] + WA 1024 20.4 24.0 21.8
DINOv2 (Vit-L) [68] 1024 56.8 67.7 62.3
DINOv2 (Vit-L) [68] + WA 1024 60.4 71.8 66.4
DINOv2 (Vit-L reg) [68] 1024 59.3 71.1 65.1
DINOv2 (Vit-L reg) [68] + WA 1024 63.9 76.6 70.6
Figure 6: Visual detection results on the High-resolution dataset using different image encoders.
@@
Dinov2 backbones with adapter. Our Weight Adapter is compatible with various backbones of Dinov2. Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements. Details of this comparison are provided in Table 7. Please refer to the Appendix for additional ablation studies.
Aggregation. For cls token embeddings, averaging the top $k$ highest (avg k) scores yields the best results [66, 53]. For FFA embeddings, the $\max$ aggregation function achieves optimal outcomes. The comparison of these aggregation functions is detailed in Table 11.
SAM variants. We evaluate different SAM variants on the High-resolution real world detection dataset. As illustrated in Table 8, SAM and our weight adapter achieves the superior performance. Given the precise bounding boxes, HQ-SAM[41] yields results comparable to those of SAM [47].
Table 11: Comparison of aggregation functions for segmentation performance. We report Average Precision (AP). avg k refers to averaging the top $k$ scores. All results are based on object proposals from Grounded SAM (GS).
Embedding Aggregation BOP Datasets
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
$cls$ token avg k 41.7 41.7 50.8 31.5 30 58 63 45.2
$cls$ token $\max$ 42 37.4 45.9 30.2 28.9 54.9 61.5 43.0
FFA avg k 42.5 42 47.4 28.2 27.3 55.1 61.5 43.4
FFA $\max$ 42.9 43 52 30.5 28.8 56.6 59.7 44.8
Runtime. We compare the efficiency of existing methods for novel instance segmentation, as presented in Table 12. Our approach significantly reduces running time by proposing only high-quality bounding boxes.
Table 12: Runtime comparisons of various methods for novel instance segmentation.
Method Proposal Server Time (sec)
CNOS FastSAM Tesla V100 0.22
CNOS DeForce RTX 3090 0.23
SAM-6D (RGB) DeForce RTX 3090 0.25
SAM-6D (RGBD) DeForce RTX 3090 0.45
CNOS SAM Tesla V100 1.84
CNOS DeForce RTX 3090 2.35
SAM-6D (RGB) DeForce RTX 3090 2.28
SAM-6D (RGBD) DeForce RTX 3090 2.80
NIDS-Net (Ours) GS RTX A5000 0.49
NIDS-Net + CA (Ours) 0.48
NIDS-Net + WA (Ours) 0.48
NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 0.48
8 Unseen Detection of BOP Benchmark
We compare our approach with ZeroPose [13], CNOS [66], and SAM-6D [53] for 2D unseen detection, as illustrated in Table 13. Our method outperforms the best RGB method by 2.5 AP and competes effectively with the top RGB-D method.
Table 13: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).
Method Proposal BOP Datasets
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
ZeroPose [13] SAM 36.7 30.0 43.1 22.8 25.0 39.8 41.6 34.1
CNOS [66] SAM 39.5 33.0 36.8 20.7 31.3 42.3 49.0 36.1
CNOS [66] FastSAM 43.3 39.5 53.4 22.6 32.5 51.7 56.8 42.8
SAM-6D (RGB) [53] FastSAM 43.8 41.7 54.6 23.4 37.4 52.3 57.3 44.4
SAM-6D (RGBD) [53] SAM 46.6 43.7 53.7 26.1 39.3 53.1 51.9 44.9
SAM-6D (RGBD) [53] FastSAM 46.3 45.8 57.3 24.5 41.9 55.1 58.9 47.1
NIDS-Net (Ours) GS 44.9 42.8 43.4 24.4 34.9 54.8 56.5 43.1
NIDS-Net + WA(Ours) 44.9 48.9 46.0 24.5 36.0 59.4 62.4 46.0
NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
9 More Qualitative Results
9.1 Adapter
To facilitate a comparison between the CLIP-Adapter and our Weight Adapter, we present a visual illustration in Figure 7. It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of background objects due to overfitting. In contrast, our Weight Adapter delivers robust embeddings within the original feature space.
9.2 Detection
We display the visual outcomes of our methodology employing the weight adapter on the LMO and YCB-V datasets in Figure 8. The gap between synthetic and real images results in some instances of detection failure. For example, some LM-O instances are not found with our method.
9.3 Segmentation
We present the visual outcomes of our approach using the weight adapter on the BoP datasets in Figures 9, 10, 11, and 12. These images demonstrate the effectiveness of our approach in cluttered environments. In some cases of T-LESS and IC-BIN datasets, Grounding DINO generates large bounding boxes which include multiple objects, causing under-segmentation. Furthermore, in IC-BIN and HB datasets, some heavily occluded objects with low confidence scores are overlooked by our method.
Figure 7: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote background objects that are erroneously classified as targets.
Figure 8: Visual detection results using the Weight Adapter on the LM-O and YCB-V datasets. Our approach detects object instances according to their synthetic template images.
Figure 9: Qualitative segmentation results on the LM-O and T-Less datasets.
Figure 10: Qualitative segmentation results on the TUD-L and IC-BIN datasets.
Figure 11: Qualitative segmentation results on the ITODD and HB datasets
Figure 12: Qualitative segmentation results on the YCB-V dataset.